OpenStudy (anonymous):

Let matrix A= |8 -2| |4 7| Let matrix B= 2A. Find b22.

OpenStudy (anonymous):

8

OpenStudy (anonymous):

sorry... its not 8, its 14

OpenStudy (mathmale):

Please try to use proper notation here. Your b22 would be clearer if written as B[2,2], which means "the element in the 2nd row, 2nd column of matrix B" Alternatively,\[B _{22}\]

OpenStudy (mathmale):

\[2\left[\begin{matrix}8 & -2 \\ 4 & 7\end{matrix}\right]=\left[\begin{matrix}16 & -4 \\ 8 & 14\end{matrix}\right]=B\]

OpenStudy (mathmale):

So, again, \[B _{22}=14\]

OpenStudy (mathmale):

Hope this helps. You might want to invest some time in learning how to type math expressions in Equation Editor (below).